# From Turbulence to Tranquility: AI-Driven Low-Altitude Network

### v1

@@ III-B ML Techniques for Interference Detection and Mitigation
Beyond spectrum sensing, ML plays an essential role in active interference management. Techniques such as reinforcement learning (RL) can allow UAVs to dynamically adjust their transmission power, channel selection, and flight trajectories to minimize interference and maximize throughput [13]. Also, generative models such as 3D-CGANs [4] can further enhance interference mitigation by predicting spatiotemporal channel occupancy maps, enabling UAVs to proactively avoid congested channels. This predictive capability is valuable in dense urban environments because the radio frequency medium can change rapidly due to UAV mobility and reflections from buildings.
Future research directions can include the development of hybrid models that integrate FL with RL to create systems that combine local adaptability with global coordination. Multi-agent RL techniques enable UAV swarms to cooperatively learn interference patterns [14]. Importantly, practical implementations need to consider the trade-offs between computational complexity, latency, and energy consumption to ensure that ML-based approaches do not overwhelm the constrained resources of aerial platforms.
@@ IV AI-Optimized Resource Allocation and Trajectory Planning
Following intelligent spectrum utilization and interference mitigation methods, efficient resource allocation and trajectory planning are a must for stable LAE. The challenge is that these tasks demand real-time decisions over multidimensional resources while having uncertain and dynamic environments. Unlike terrestrial systems, since LAE networks are characterized by high mobility, strict energy and delay constraints, a holistic design approach jointly considering UAV trajectory control and communication resource scheduling is demanded. Hence, this section discusses the possible solutions regarding joint resource allocation and trajectory planning.
@@ IV-A Reinforcement Learning–Based Resource Management in Dynamic Airspace
For example, in [15], it is shown that the digital twin network (DNT) combining gated recurrent units (GRUs) with value decomposition networks (VDN) enables distributed agents like UAVs and base stations to collaboratively optimize spectrum and computing resources. This approach outperforms independent Q-learning baselines, showing the power of cooperative learning. Another interesting study [13] combines apprenticeship learning and inverse reinforcement learning (IRL) to create optimal policies from expert demonstrations. IRL models capture the intent behind expert actions, which allows UAVs to generalize better than traditional behavioral cloning. This is especially useful in LAE environments where manual design of reward functions is difficult due to complex, dynamic objectives.
@@ IV-B Optimization of UAV Trajectories to Minimize Interference
Moreover, trajectory optimization can improve fairness and quality of service by balancing coverage across users and reducing interference. Future research should investigate the integration of online learning and transfer learning to enable aerial nodes to generalize trajectory strategies across missions and environments.
@@ V Experimental Validation and Testbed Deployment
Besides the AI-driven frameworks and optimization strategies discussed in the previous sections, real-world validation is essential. Experimental validation not only demonstrates the feasibility of theoretical models but also bridges the gap between simulation and practical deployment and provides critical insights for practical LAE networks.
Figure 3: The AERPAW platform can operate in three experimental modes: physical testbed, sandbox, and digital twin environment.
@@ V-A AERPAW: An Experimental Platform for LAE Research
AERPAW, located in Raleigh, NC, serves as a flagship testbed within the NSF Platforms for Advanced Wireless Research (PAWR) program and is uniquely tailored for LAE experimentation. AERPAW, consisting of programmable UAVs and ground nodes equipped with SDRs, offers a highly flexible platform for controlled wireless experiments in heterogeneous environments [8]. By spanning both urban and rural areas (e.g., Lake Wheeler Field Labs and the NC State Centennial Campus), AERPAW enables reproducible and scalable experiments for spectrum management strategies and trajectory optimization techniques discussed above.
As AERPAW combines live outdoor experimentation and digital twin–based emulation, it provides a powerful framework for researchers to design, simulate, and validate aerial wireless communication scenarios before field deployment. This capability is critical to validate the methods and architectures towards regulation, standardization, and practical usage since it allows for iterative refinement of LAE systems.
@@
V-B Performance Metrics for LAE Validation
Validating LAE technologies at AERPAW involves a comprehensive set of metrics aligned with the core challenges discussed in Section II. These metrics include but are not limited to signal quality, link availability, outage probability, end-to-end throughput, channel rank for multi-antenna systems, latency, and spectrum occupancy. The testbed’s support for packet-level, IQ-level, and RF-level emulation enables fine-grained analysis of such key performance indicators (KPIs) across a wide range of deployment conditions.
Altitude-dependent measurements of propagation conditions, spectrum occupancy, and other KPIs, especially in sub-6 GHz bands, provide valuable insights into the spatial and temporal dynamics of LAE networks, as well as the suitability of specific spectrum bands for supporting different LAE use cases. Experiments such as those conducted at AERPAW are critical, as traditional simulation tools and theoretical analysis often fail to realistically characterize the KPIs of real-world LAE deployments.
V-C Deployment Insights and Testbed Lessons
@@
Practical experience with AERPAW yields several important lessons for LAE development. First, trajectory-aware communication strategies are indispensable under the changing propagation conditions associated with UAV mobility, altitude, and orientation. Second, bridging the simulation-to-reality gap remains a nontrivial challenge. Models trained in digital twin environments often require calibration with real-world measurements to achieve reliable operation. Finally, the modular design of AERPAW’s portable radio nodes and its abstraction layers enables flexible, future-proof architectures capable of accommodating heterogeneous hardware and regulatory requirements.
@@
Fig. 3 presents representative results that evaluate the occupancy of altitude-dependent spectrum in rural and urban environments for LAE networks. In particular, a helikite is used to collect spectrum occupancy data at various altitudes, locations, and years, between 85 MHz to 6 GHz. The cumulative distribution function (CDF) of average received power for LTE downlink Band 13, grouped by UAV altitude intervals, is provided. At lower altitudes (0–40 m), aggregate signal power measurements in Band 13 are significantly weaker, with most measurements falling below -40 dBm. As altitude increases, especially beyond 80 m, received power improves considerably due to enhanced LOS conditions, with the majority of values exceeding -20 dBm.
Also in Fig. 3, we show heatmaps of spectrum occupancy across multiple frequency bands of relevance to LAE networks, for both urban and rural environments, comparing conditions below and above 50 m altitude. We consider the spectrum to be occupied if the received signal power for a specific altitude and frequency exceeds a set threshold. At lower altitudes, spectrum occupancy exhibits significant variation. In earlier urban measurements, some bands (e.g., ISM) show minimal activity, while others (e.g., Bands 13 and 14) are heavily utilized. At higher altitudes, spectrum usage is consistently denser across almost all bands and years, especially in n41 and ISM bands. These findings highlight the importance of altitude-aware spectrum analysis for effective LAE network planning to support different use cases.
@@
The unique characteristics and statistics of the real-world data, such as for altitude-dependent spectrum occupancy, underscore the importance of AI-driven solutions presented in this study. Such data-driven AI solutions can be used to improve spectrum sensing, interference mitigation, resource allocation, and trajectory optimization for LAE networks. They also highlight the need for integrated C2S frameworks and standardized testbed practices. AERPAW plays a critical role in data generation and standardization efforts for LAE networks. The testbed can serve as a validation platform for emerging protocols related to spectrum coexistence, trajectory planning, and resource optimization. By supporting standardized interfaces and benchmarks, AERPAW can help ensure that AI-enhanced solutions are not only technically feasible but also interoperable and aligned with evolving global standards.
@@ VI Open Issues
Moving forward, several open research directions and technical challenges must be addressed to fully realize the potential of AI-driven LAE networks. The primary concern is the integration of explainable AI or trustworthiness techniques into decision-making frameworks. As AI models impact mission-critical operations, it becomes important to ensure transparency and interpretability in their decisions, especially under uncertain or adversarial conditions.
@@
Furthermore, establishing interoperable standards between air, space, and terrestrial systems remains a challenge. Cooperation with satellite systems has been studied in the literature, but more efforts are needed for seamless integration that requires well-defined protocols and spectrum policies.
@@ VII Concluding Remarks
The emergence of LAE networks presents an exciting frontier in wireless communications as offering transformative capabilities across several domains such as aerial transportation, environmental monitoring, and logistics. However, realizing these promising capabilities requires overcoming challenges in communication, sensing, and system coordination within highly dynamic and resource-constrained airspace environments. In this study, we provide a comprehensive exploration of the key subsystems required to support the vision of AI-driven LAE networks. Along with detailing the challenges towards full operable LAE networks, we identify and discuss three critical issues as follows: spectrum sensing and coexistence strategies empowered by ML, RL-driven resource allocation and trajectory planning, and the standardization and experimental validation of proposed solutions through real-world testbeds such as AERPAW.
We indicate that ML-based approaches can provide significant promise in adapting to the non-stationary, heterogeneous conditions of LAE. For instance, FL models, particularly those incorporating SNR-aware aggregation schemes, can support decentralized and privacy-preserving spectrum sensing while maintaining high accuracy in evolving environments.
We also highlight the role of experimental platforms like AERPAW in narrowing the gap between theoretical innovation and practical deployment. Through digital twins, sandbox environments, and live testbeds, AERPAW enables validation of AI-supported LAE solutions under realistic conditions. These efforts are essential not only for technical development but also for ensuring regulatory compliance, interoperability, and robustness across diverse deployment scenarios.
@@
KÜRŞ̧AT TEKBIYIK [StM’19, M’25] (ktekbiyik@gmail.com) received his Ph.D. degree in telecommunications engineering from Istanbul Technical University, Istanbul, Türkiye. His research interests include machine learning applications in wireless communications, terahertz communications, and non-terrestrial networks.
@@
ANTOINE LESAGE-LANDRY [StM’16, M’20, SM’24] is an Associate Professor in the Department of Electrical Engineering at Polytechnique Montréal, QC, Canada. His research focuses on decision-making in complex engineering systems like electric power systems and wireless communication networks.

### v2

@@ III-B ML Techniques for Interference Detection and Mitigation
Beyond spectrum sensing, ML plays an essential role in active interference management. Techniques such as reinforcement learning (RL) can allow UAVs to dynamically adjust their transmission power, channel selection, and flight trajectories to minimize interference and maximize throughput. Also, generative models such as 3D-CGANs [8] can further enhance interference mitigation by predicting spatiotemporal channel occupancy maps, enabling UAVs to proactively avoid congested channels. This predictive capability is valuable in dense urban environments because the radio frequency medium can change rapidly due to UAV mobility and reflections from buildings.
Future research directions can include the development of hybrid models that integrate FL with RL to create systems that combine local adaptability with global coordination. Multi-agent RL techniques enable UAV swarms to cooperatively learn interference patterns [14]. In real applications, it is important to balance computing power, delay, and energy use so that ML methods do not exceed the limited resources of aerial vehicles unlike the current literature trying to optimize only one of them.
@@ IV AI-Optimized Resource Allocation and Trajectory Planning
Following intelligent spectrum utilization and interference mitigation methods, efficient resource allocation and trajectory planning are a must for stable LAE. The challenge is that these tasks demand real-time decisions over multidimensional resources while having uncertain and dynamic environments. Unlike terrestrial systems, because LAE networks are characterized by high mobility, strict energy and delay constraints, a holistic design approach jointly considering UAV trajectory control and communication resource scheduling is demanded. Hence, this section discusses the possible solutions regarding joint resource allocation and trajectory planning.
@@ IV-A Reinforcement Learning–based Resource Management in Dynamic Airspace
For example, in [15], it is shown that the digital twin network (DNT) combining gated recurrent units (GRUs) with value decomposition networks (VDN) enables distributed agents like UAVs and base stations to collaboratively optimize spectrum and computing resources. This approach outperforms independent Q-learning baselines, showing the power of cooperative learning. This is especially useful in LAE environments where manual design of reward functions is difficult due to complex, dynamic objectives.
@@ IV-B Optimization of UAV Trajectories to Minimize Interference
Moreover, trajectory optimization can improve fairness and quality of service by balancing coverage across users and reducing interference [2]. However, many works assume idealized channel and interference models or simple network topologies which may not fully capture dynamic mobility and heterogeneous nodes. Another important point missed by these works is nonideal flight dynamics consisting of no‐fly zones, regulatory and safety constraints, 3D mobility, weather, changes in ground network loads. Thus, bridging to the real world is still a challenge.
@@ V Experimental Validation and Testbed Deployment
Besides the AI-driven frameworks and optimization strategies discussed in the previous sections, real-world validation is essential. Experimental validation not only demonstrates the feasibility of theoretical models but also bridges the gap between simulation and practical deployment and provides critical insights for practical LAE networks. In this section, we introduce AERPAW along with performance metrics regarding the validation of LAE systems and some practical insights learned from the testbed for field-deployments.
Figure 4: The AERPAW platform supports three operational modes: physical testbed, sandbox, and digital twin environments. In the CDF plot, solid and dashed curves correspond to measurements taken below and above 50 m altitude, respectively. The vertical black dotted line represents the threshold value of 45-45 dBm, which is used to determine the spectrum occupancy across frequency bands.
@@ V-A AERPAW: An Experimental Platform for LAE Research
AERPAW, located in Raleigh, NC, serves as a flagship testbed within the NSF Platforms for Advanced Wireless Research (PAWR) program and is uniquely tailored for LAE experimentation. AERPAW, consisting of programmable UAVs and ground nodes equipped with SDRs, offers a highly flexible platform for controlled wireless experiments in heterogeneous environments [7]. By spanning both urban and rural areas e.g., Lake Wheeler Field Labs and the NC State Centennial Campus, AERPAW enables reproducible and scalable experiments for spectrum management strategies and trajectory optimization techniques discussed above.
As AERPAW combines live outdoor experimentation and digital twin–based emulation, it provides a powerful framework for researchers to design, simulate, and validate aerial wireless communication scenarios before field deployment. This capability is critical to validate the methods and architectures towards regulation, standardization, and practical usage because it allows for iterative refinement of LAE systems.
@@
V-B Deployment Insights and Testbed Lessons
@@
The practical experience with AERPAW yields several important lessons for LAE development. First, trajectory-aware communication strategies are indispensable under the changing propagation conditions associated with UAV mobility, altitude, and orientation. Second, bridging the simulation-to-reality gap remains a nontrivial challenge. Models trained in digital twin environments often require calibration with real-world measurements to achieve reliable operation. Finally, the modular design of AERPAW’s portable radio nodes and its abstraction layers enables flexible, future-proof architectures capable of accommodating heterogeneous hardware and regulatory requirements.
@@
Fig. 4 presents representative results illustrating altitude-dependent spectrum characteristics in rural and urban environments for LAE networks. A helium-assisted kite (helikite) was employed to collect wideband spectrum measurements between 85 MHz and 6 GHz at multiple altitudes, locations, and measurement periods. The empirical cumulative distribution function (CDF) of the average received power from the Urban 2025 measurement campaign is presented for LTE downlink Band 12, LTE uplink Band 12, and 5G Band n41, with results grouped by UAV altitudes below and above 50 m. The CDFs reveal that at lower altitudes, received signal power is generally weaker, whereas higher altitudes yield noticeably stronger levels due to improved LoS conditions, with most samples exceeding the 45-45 dBm threshold.
The same figure includes heatmaps depicting spectrum occupancy across several frequency bands of relevance to LAE networks, comparing urban and rural environments below and above 50 m altitude. For each band and altitude group, occupancy is determined from the mean received power across the corresponding band, which is considered occupied when this mean exceeds 45-45 dBm. At lower altitudes, occupancy patterns vary considerably; for instance, LTE downlink Band 12 shows limited activity in Urban 2022 but significant utilization in Rural 2024. At higher altitudes, spectrum occupancy becomes denser and more uniform across most bands and years. These results confirm that both signal availability and interference conditions are strongly influenced by altitude, emphasizing the importance of altitude-aware spectrum characterization for reliable and efficient LAE network deployment.
The unique characteristics and statistics of the real-world data, such as for altitude-dependent spectrum occupancy, underscore the importance of AI-driven solutions presented in this study. Such data-driven AI solutions can be used to improve spectrum sensing, interference mitigation, resource allocation, and trajectory optimization for LAE networks. They also highlight the need for integrated C2S frameworks and standardized testbed practices. Specifically, these results bridge the gap between the theoretical C2S challenges introduced in Section II-B and practical deployment. For instance, the altitude-dependent spectrum heterogeneity shown in the heatmaps implies that the radio environment is too dynamic, and therefore, static rule-based models could struggle. A collaborative C2S framework must integrate such real-world sensing capabilities to allocate resources. These experiments validate the need for a holistic C2S design for LAE networks because the results illustrate that physical sensing constrains AI models.
Moreover, these measurements provide real-world data required to train and validate the AI models discussed in the previous sections. For example, the altitude-dependent path loss and interference data collected from the testbed are used to train Physics-Inspired Kolmogorov-Arnold Networks (PIKAN) [11]. PIKAN leverages this real-world dataset to learn symbolic, interpretable expressions of the channel that also employ physical propagation laws. This demonstrates that the AERPAW setups act as an enabler for AI model training and validation to capture complex non-stationary airspace behaviors such as elevation and altitude-dependent fading, while theoretical simulations often oversimplify these impacts.
@@
AERPAW plays a critical role in data generation, model validation, and standardization efforts for LAE networks. The testbed can serve as a validation platform for emerging protocols related to spectrum coexistence, trajectory planning, and resource optimization. By supporting standardized interfaces and benchmarks, AERPAW can help ensure that AI-enhanced solutions are not only technically feasible but also interoperable and aligned with evolving global standards.
@@ VI Open Issues
Although we mention some open issues specific to research areas given above, this section aims to discuss open issues from a broader perspective, without narrowing to specific tasks such as spectrum sensing. We aim to provide holistic future directions toward AI-driven LAE.
As LAE networks integrate into critical infrastructure, the security of their AI-driven components needs to be a crucial concern. As we explained, distributed frameworks like FL can reduce data leakage, but model poisoning and adversarial attacks might manipulate the global model. In spectrum sensing or trajectory planning, a malicious edge node could change training weights to mask interference or induce collisions. Consequently, deploying AI in LAE requires sophisticated aggregation protocols and anomaly detection mechanisms so that these systems can improve safety and maintain operational integrity. However, the primary concern is the integration of explainable AI and trustworthiness techniques into decision-making frameworks. As AI models impact mission-critical operations, it becomes important to ensure transparency and interpretability in their decisions, especially under uncertain or adversarial conditions. To address this, recent frameworks such as PIKAN [11] is proposed to bridge the gap between DL accuracy and physical interpretability. Drawing inspiration from physical principles such as path loss exponents and two-ray models, PIKAN learns channel behaviour and provides symbolic expressions that explain the underlying logic of the predictions. Such physics-inspired and explainable approaches are required for certification and regulatory approval because aviation authorities audit and trust the decisions made by autonomous aerial systems based on those.
@@
Besides, establishing interoperable standards between air, space, and terrestrial systems remains a challenge. Cooperation with satellite systems has been studied in the literature, but more efforts are needed for seamless integration that requires well-defined protocols and spectrum policies.
@@ VII Concluding Remarks
The emergence of LAE networks presents an exciting frontier in wireless communications as offering transformative capabilities across several domains such as aerial transportation, environmental monitoring, and logistics. However, realizing these promising capabilities requires overcoming challenges in communication, sensing, and system coordination within highly dynamic and resource-constrained airspace environments. In this study, we provide a comprehensive exploration of the key wireless communication subsystems required to support the vision of AI-driven LAE networks. Along with detailing the challenges towards full operable LAE networks, we identify and discuss three critical issues as follows: spectrum sensing and coexistence strategies empowered by ML, RL-driven resource allocation and trajectory planning, and the standardization and experimental validation of proposed solutions through real-world testbeds.
ML-based approaches can provide promise in adapting to the non-stationary, heterogeneous conditions of LAE. For instance, FL models, particularly those incorporating SNR-aware aggregation schemes, can support decentralized and privacy-preserving spectrum sensing while maintaining high accuracy in evolving environments. It should be noted that there are several open issues for practical deployment. We also highlight the role of experimental platforms like AERPAW in narrowing the gap between theoretical innovations and practical deployments. Through digital twins, sandbox environments, and live testbeds, they can enable validation of AI-supported LAE solutions under realistic conditions. These efforts are essential not only for technical development but also for ensuring regulatory compliance, interoperability, and robustness across diverse deployment scenarios.
@@
KÜRŞ̧AT TEKBIYIK [StM’19, M’24] (kursat.tekbiyik@polymtl.ca) received his Ph.D. degree in telecommunications engineering from Istanbul Technical University, Istanbul, Türkiye. His research interests include machine learning applications in wireless communications, terahertz communications, and non-terrestrial networks.
@@
ANTOINE LESAGE-LANDRY [StM’16, M’20, SM’24] is an Associate Professor of Electrical Engineering at Polytechnique Montréal, QC, Canada. His research focuses on decision-making in complex engineering systems like electric power systems and wireless communication networks.